BACCK Inc. currently has an EPS of $5.23 and an earnings growth rate of 5 percent. If the benchmark PE ratio is 14, what is the target price six years from now? A. $77 B. $89 C. $98 D. $81
Answer #1

Expected EPS=5.23*(1+Growth Rate)^6

=5.23*(1.05)^6

=$7.0087002

PE ratio=Stock price/EPS

Hence expected stock price=7.0087002*14

=$98(Approx).